Find an equation of the line having the given slope and containg the given point.
m=2/3 ,(2,-9)
The equation of the line is y=

Answer by fcabanski(1391)

Whenever you're given a point and a slope, you find the equation of the line by:


Step 1 plug the point's x and y coordinates, and the slope, into the general equation y=mx+b.


y=mx+b becomes -9=(2/3)2 + b ---> -9 = 4/3 +b (subtract 4/3 from both sides):


-9 - 4/3 = b ---> -(27/3) -(4/3) = b = -(31/3)


Step 2 rewrite the general formula y=mx+b with the b found in step one, and the given slope (m).


y=(2/3)x-(31/3)
Check my profile and contact me for one on one online tutoring.